OpenStudy (thatonegirl_):

find the inverse: f(x)=1/2+7^x ... please explain >.<

OpenStudy (anonymous):

step 1: \[f(x)-1/2=7^{x}\] step 2: \[\log(f(x)-1/2) = \log{7^{x}} \] step 3: \[\log(f(x)-1/2)=xlog7\] step4: \[[\log(f(x) - 1/2)]/\log7]=x\] step 5: \[[logx-1/2]/\log7=f(x)\]

OpenStudy (thatonegirl_):

Thanks! :)